OpenStudy (anonymous):

solvethe following sytem of equations... show solution method 3x+2y=5,-2x+y=-1

OpenStudy (amistre64):

What to do, what to do... We can eliminate a variable; lets pick 'y' just cause.. 3x +2y = 5 (-2)(-2x + y = -1) 3x +2y = 5 + 4x -2y = 2 ------------ 7x + 0y = 7 x = 1 if I did it right :) now pick an equation and substitute x=1 and solve for y: 3(1) +2y = 5 2y = 2 y=1 that was nifty; but lets see if these values for x and y work with the other equation as well, otherwise it wont be a solutions. ,-2(1)+(1)=-1 -2 +1 = -1 i think its true :)